# Reconstructing the electro-weak theory from a gauge theory

### v1

@@ Abstract
$U(4)$ local transformations on the four Weyl spinors forming the isospin doublet of Dirac fermions are assumed as symmetries of the standard model. With the Lorentz transformations considered simultaneously, the symmetry group is enlarged in order to form a closed Lie algebra. In this framework, the chirality mixing gauge components with certain constraints collectively are identified as the Higgs field in the standard model. The scalar-appearance of the gauge-natured Higgs field and its varying coupling constants with the fermions, i.e. the mass-ratio parameters, are also given natural explanation. Additionally, a background of a constant right-handed gauge component is postulated to obtain the symmetry-breaking Higgs potential, which leads to a possible explanation of the parity violation in weak interactions. Further, the new framework provides an alternative perspective to understand the hyper symmetry $U(1)_{Y}$, as a combination of a simultaneous phase rotation on both the left and right handed fermions and a rotation in the third direction of $SU(2)_{R}$ on the right-handed fermions. At last, a topological term involving the Higgs field is found to exist.
@@ I Introduction
The standard model of particle physics is by far the most successful theory. The development of the theory has gone through a series of conceptual and theoretical breakthroughs, such as the symmetry-breaking concept introduced to particle physics by Nambu, the Higgs mechanism innovated by Higgs, Englert et.al, the proof of the renormalisability of Yang-Mills theories by ’t Hooft and Veltman, and the electro-weak unification model invented by Glashow, Weinberg and Salam, only to name the most influential. Physicists’ confidence in the standard model was consolidated in 2012 when the predicted Higgs particle was eventually discovered. The extensive experimental research before and after the discovery also accumulated even more faith in the standard model. Despite the great success, there are still some concerns about the CP violation, the value of muon $g-2$ and the Higgs-mass problem. These concerns of course do not tarnish the glory of the standard model at all. But the desire for removing every single piece of doubt encourages physicists not to cease in searching for new physics.
@@ II.1 The Lagrangian and the symmetry transformations
and where now $\bar{\Psi}=\Psi^{\dagger}\Gamma^{0}$, ∂̸$\not{\partial}=\Gamma^{\mu}\partial_{\mu}∂$ with $\Gamma^{0}=\sigma^{1}\otimes I_{2}\otimes I_{2}$ and $\Gamma^{k}=i\sigma^{2}\otimes I_{2}\otimes\sigma^{k}.$ Further $\Gamma^{0}\Gamma^{0}=I_{8}$ and $\Gamma^{0}\Gamma^{k}=-\sigma^{3}\otimes I_{2}\otimes\sigma^{k}.$ Let’s assume that this Lagrangian admits the symmetry transformations of $U(4)$ among the four massless Weyl fermions:
@@ IV.1.1 The gauge configuration for deriving the Standard-Model Higgs field
Under the Lorentz transformation $\Lambda_{\frac{1}{2}m}=\exp(\theta\sigma^{3}\otimes I_{2}\otimes\sigma^{*})\cdot\exp(i\alpha I_{2}\otimes I_{2}\otimes\sigma^{1})$ with $\sigma^{*}=\cos 2\alpha\sigma^{2}-\sin 2\alpha\sigma^{3}$ $A_{mix,\mu}$ are transformed into
$\begin{split}A_{mix,0}\Rightarrow&\cosh 2\theta\phi(x)\sigma^{1}\otimes I_{2}\otimes I_{2}\\ &-i\sinh 2\theta\phi(x)\sigma^{2}\otimes I_{2}\otimes\sigma^{*},\\ A_{mix,1}\Rightarrow&-i\phi(x)\sigma^{2}\otimes I_{2}\otimes\sigma^{1},\\ A_{mix,2}\Rightarrow&i\cosh 2\theta\phi(x)\sigma^{2}\otimes\sigma^{3}\otimes\sigma^{*}\\ &-\sinh 2\theta\phi(x)\sigma^{1}\otimes\sigma^{3}\otimes I_{2},\\ A_{mix,3}\Rightarrow&-i\cos 2\alpha\phi(x)\sigma^{2}\otimes\sigma^{3}\otimes\sigma^{3}\\ &-i\sin 2\alpha\sigma^{2}\otimes\sigma^{3}\otimes\sigma^{2}.\end{split}$ (39)
@@
with $m_{a}+m_{b}=4\sinh^{2}\theta\cos^{2}\alpha$ and $m_{a}-m_{b}=4\sinh^{2}\theta\sin^{2}\alpha$ the mass ratios of fermions in each generation.
@@ IV.1.3 The Lagrangian of the electroweak theory—derived from a pure gauge theory
$\begin{split}L_{t}=&\sum_{m}\Psi_{m}^{\dagger}h\Gamma^{\mu}(i\partial_{\mu}+A_{L\mu}+B_{\mu})\Psi_{m}\\ &+\sum_{m}\Psi_{m}^{\dagger}\Phi_{m}\Psi_{m}\\ &+\frac{1}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi-\frac{g^{2}}{3}[(\Phi^{\dagger}\Phi)^{2}-2\tilde{v}^{2}\Phi^{\dagger}\Phi]\\ &-\frac{1}{4}\mathbf{tr}(F_{L\mu\nu}F_{L}^{\mu\nu}+f_{Y\mu\nu}f_{Y}^{\mu\nu})\\ &+3g^{2}(A_{R\mu}^{3}-A_{L\mu}^{3})v^{\mu}\end{split}$ (50)
@@
with $\bar{\Phi}=i\sigma^{2}\Phi^{*}.$ The details of the calculation are presented in the appendix. An immediate observation is that there is an extra-term $3g^{2}v_{\mu}(A^{3}_{L\mu}-A^{3}_{R\mu})$. We have not performed analysis on the quantized theory. Hopefully there are some processes that can help determine the values of $v_{\mu}.$
@@ IV.2.1 Deriving the abelian Higgs model from a gauge theory
Now just as in the non-abelian case when an internal Lorentz transformation is performed with $\Lambda_{\frac{1}{2}m}=\exp(\theta\sigma^{3}\otimes\sigma^{*})\exp(i\alpha I_{2}\otimes\sigma^{1})$ with $\sigma^{*}=\cos 2\alpha\sigma^{2}-\sin 2\alpha\sigma^{3}$, these chirality mixing gauge components are transformed into:
$\begin{split}A_{mix,0}&\Rightarrow\cosh 2\theta\phi(x)\sigma^{1}\otimes I_{2}-i\sinh 2\theta\phi(x)\sigma^{2}\otimes\sigma^{*}\\ A_{mix,1}&\Rightarrow i\phi(x)\sigma^{2}\otimes\sigma^{1}\\ A_{mix,2}&\Rightarrow i\cosh 2\theta\phi(x)\sigma^{2}\otimes\sigma^{*}-\sinh 2\theta\phi(x)\sigma^{1}\otimes I_{2}\\ A_{mix,3}&\Rightarrow i\cos 2\alpha\phi(x)\sigma^{2}\otimes\sigma^{3}+i\sin 2\alpha\sigma^{2}\otimes\sigma^{2}.\end{split}$ (58)
@@
which results in a mass ratio parameter $m=4\cosh^{2}\theta\cos^{2}\alpha.$
@@
$\begin{split}-\frac{1}{8}\mathbf{tr}ZF_{\mu\nu}ZF^{\mu\nu}+h.c.=&-\frac{1}{4}f_{L\mu\nu}f_{L}^{\mu\nu}-\frac{1}{4}f_{R\mu\nu}f_{R}^{\mu\nu}\\ &+\frac{1}{2}(D_{\mu}\phi)^{\dagger}D^{\mu}\phi\\ &-\frac{e^{2}}{3}(\varphi^{\dagger}\varphi)^{2}+\frac{e^{2}}{2}v_{\mu}v^{\mu}\varphi^{\dagger}\varphi\\ &+3e^{2}(b_{\mu}^{3}-a_{L\mu}^{3})v^{\mu}\end{split}$ (60)
@@ V Problems and further research directions
At last, the ”over-size” of the symmetry groups in both cases raises serious concerns, as in reality we have observed only a very small portion of the interactions described by the two symmetry groups. If some of symmetries were not to be gauged, what would be the guiding principle about which to be gauged and which not? The geometric-algebraic perspective may leads us to realise a possibility to shrink the size of the two groups. In a geometric algebra, the bi-vectors can be naturally interpreted as generators of rotations in the geometric space. In either of the abelian and non-abelian cases, mapping the complexified geometric algebra to a real one with one more dimension, then the bi-vectors can be selected to generate a smaller group with obvious geometric meaning. Then mapping these bi-vectors back to the complexfied algebra gives a Lie-subgroup of the original group. In the non-abelian case this would give a Lie-subgroup generated by 21 generators, including the six Lorentz generators. In the abelian case this would give a Lie-subgroup generated by 10 generators, including the six Lorentz generators. However, the problem is that some of the desired generators, for instance, the chirality mixing generators associated to the Higgs field, are not included in the Lie-algebra of these Lie subgroups. Another direction to look at is the transformation of the spinor metric $h.$ Under some symmetry transformations the spinor metric does not stay invariant. If the invariance of $h$ is set as a criteria for whether or not to gauge the symmetry, transformations generated by $\sigma^{L}\otimes\sigma^{k}\otimes I_{2}$ would be excluded from the gauged symmetries. This cannot be true. In fact, $\sigma^{L}\otimes\sigma^{k}\otimes I_{2}$ commutes with the products $h\Gamma^{\mu}$ which guarantees the products as whole entities invariant. But the products $h\Gamma^{\mu}$ would be altered by a Lorentz transformation. These arguments suggest that a simple criteria related to $h$ cannot be working. For a working criterion to be found, a clear understanding of the nature of the spinor metric is probably the key.
@@ VI Appendices
leading to the expression of the $\Gamma^{\mu}s$ given in Eq. (LABEL:eq:_Gamma). The Pauli matrices are:
@@ VI.2 Appendix B— Calculation of the Higgs sector of the Lagrangian
$\begin{split}-\frac{1}{8}&\mathbf{tr}ZF_{\mu\nu}ZF^{\mu\nu}+h.c.=-\frac{1}{4}\{\mathbf{tr}F_{L\mu\nu}F_{L}^{\mu\nu}+\mathbf{tr}f_{R\mu\nu}f_{R}^{\mu\nu}\\ &+\mathbf{tr}f_{N\mu\nu}f_{N}^{\mu\nu}-\mathbf{tr}(D_{\mu}\eta_{\nu}-D_{\nu}\eta_{\mu})(D^{\mu}\eta^{\nu}-D^{\nu}\eta^{\mu})\\ &-g^{2}\mathbf{tr}[\eta_{\mu},\eta_{\nu}][\eta^{\mu},\eta^{\nu}]\\ &+g^{2}\mathbf{tr}([C_{3R\mu},\eta_{\nu}]-[C_{R\nu},\eta_{\mu}])([C_{3R}^{\mu},\eta^{\nu}]-[C_{R}^{\nu},\eta^{\mu}])\\ &-2ig\mathbf{tr}(F_{L\mu\nu}+f_{R\mu\nu}+f_{N\mu\nu})[\eta^{\mu},\eta^{\nu}]\\ &+2ig\mathbf{tr}(D_{\mu}\eta_{\nu}-D_{\nu}\eta_{\mu})([C_{3R}^{\mu},\eta^{\nu}]-[C_{R}^{\nu},\eta^{\mu}])\}\\ &=-\frac{1}{4}\mathbf{tr}F_{L\mu\nu}F_{L}^{\mu\nu}-\frac{1}{4}\mathbf{tr}f_{R\mu\nu}f_{R}^{\mu\nu}-\frac{1}{4}\mathbf{tr}f_{N\mu\nu}f_{N}^{\mu\nu}\\ &+\frac{3}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi-3g^{2}(\Phi^{\dagger}\Phi)^{2}+\frac{3}{2}g^{2}v_{\mu}v^{\mu}\Phi^{\dagger}\Phi\\ &+3g^{2}(A_{R\mu}^{3}-A_{L\mu}^{3})v^{\mu}\end{split}$ (75)
@@
Calculating the relevant terms in $-\frac{1}{4}ZF_{\mu\nu}ZF^{\mu\nu}$ gives the Lagrangian of the Higgs sector
$L_{\Phi}=\frac{1}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi-\frac{g^{2}}{3}[(\Phi^{\dagger}\Phi)^{2}-2\tilde{v}^{2}\Phi^{\dagger}\Phi]$ (77)
If the right-handed gauge field components $A_{R\mu}^{1,2}\sigma^{R}\sigma^{1,2}\otimes I_{2}$ are restored in the theory, the term $\frac{g^{2}}{4}\mathbf{tr}([C_{3R\mu},A_{R\nu}]-[C_{3R\nu},A_{R\mu}])([C_{3R}^{\mu},A_{R}^{\nu}]-[C_{3R}^{\nu},A_{R}^{\mu}])$ would produce two terms $-g^{2}v_{\mu}v^{\mu}A^{1}_{R\nu}A_{R}^{1\nu}+g^{2}v_{\mu}v^{\nu}A^{1}_{R\nu}A^{1\mu}_{R}$ and $-g^{2}v_{\mu}v^{\mu}A^{2}_{R\nu}A_{R}^{2\nu}+g^{2}v_{\mu}v^{\nu}A^{2}_{R\nu}A^{2\mu}_{R}$ where in the summation $\mu\neq\nu.$ When the gauge $A_{R0}^{1,2}=0$ is chosen and in the frame where $v_{k}=0$ this gives $+g^{2}v_{0}v^{0}A^{1}_{k}A^{1}_{k}$ and $+g^{2}v_{0}v^{0}A^{2}_{k}A^{2}_{k}$. Given that the vector $v_{\mu}$ must be timelike, the sign of these two terms seems to be opposite to that of a bosonic mass term in a Lagrangian. However, these terms are essentially the $E_{k}E^{k}$ part of electric-magnetic Lagrangian. Thus in the Hamiltonian, they would not change the sign and forming mass terms with correct signs. Thus, these two terms together with $\frac{1}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi$ produce six mass terms with the least mass mass values,
@@
While the ${W^{\pm}}$ field components would acquire a mass coefficient $\frac{1}{2}g^{2}\tilde{v}^{2}$ from $\frac{1}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi$, it can be predicted that these right-handed weak gauge bosons have masses at least $1.9$ times of the left-handed $W^{\pm}$ bosons, a bit too heavy to be discovered at the currently running colliders. Also they are too heavy to be excited which results in the disparity between the left-handed and right-handed weak interactions.
@@
To reconstruct the abelian Higgs model, let us assume that the gauge field takes up the following configuration: $A_{\mu}=A_{L\mu}+A_{R\mu}+C_{R\mu}+\eta_{\mu}$ with $C_{R\mu}=v_{\mu}\sigma^{R}\otimes I_{2}$ (actually $C_{L\mu}=v_{\mu}\sigma^{L}\otimes I_{2}$ and $C_{\mu}=v_{\mu}\sigma^{3}\otimes I_{2}$ equally do the job) and
@@
$\begin{split}-\frac{1}{8}\mathbf{tr}ZF_{\mu\nu}ZF^{\mu\nu}+h.c.=&-\frac{1}{4}\mathbf{tr}f_{L\mu\nu}f_{L}^{\mu\nu}-\frac{1}{4}\mathbf{tr}f_{R\mu\nu}f_{R}^{\mu\nu}\\ &+\frac{1}{2}(D_{\mu}\varphi)^{\dagger}D^{\mu}\varphi\\ &-\frac{e^{2}}{3}(\varphi^{\dagger}\varphi)^{2}+\frac{e^{2}}{2}v_{\mu}v^{\mu}\varphi^{\dagger}\varphi\\ &+3g^{2}(b_{\mu}^{3}-a_{L\mu}^{3})v^{\mu}\end{split}$ (80)

### v2

# Reconstructing the electro-weak theory from a pure gauge theory

@@ Abstract
$U(4)$ local transformations on the four Weyl spinors forming the isospin doublet of Dirac fermions are assumed as symmetries of the standard model. With the Lorentz transformations considered simultaneously, the symmetry group is enlarged in order to form a closed Lie algebra. In this framework, the chirality mixing gauge components with certain constraints collectively are identified as the Higgs field in the standard model. The scalar-appearance of the gauge-natured Higgs field and its varying coupling constants with the fermions, i.e. the mass-ratio parameters, are also given natural explanation. Additionally, a background of a constant right-handed gauge component is postulated to obtain the symmetry-breaking Higgs potential, which also leads to a possible explanation to the parity violation in weak interactions. Further, the new framework provides an alternative perspective to understand the hyper symmetry $U(1)_{Y}$, as a combination of a simultaneous phase rotation on both the left and right handed fermions and a rotation in the third direction of $SU(2)_{R}$ on the right-handed fermions. At last, a topological term involving the Higgs field is found to exist.
@@ I Introduction
The standard model of particle physics is by far the most successful theory. The development of the theory has gone through a series of conceptual and theoretical breakthroughs, such as the symmetry-breaking concept introduced to particle physics by Nambu, the Higgs mechanism innovated by Higgs, Englert et.al, the proof of the renormalisability of Yang-Mills theories by ’t Hooft and Veltman, and the electro-weak unification model invented by Glashow, Weinberg and Salam, only to mention the most influential and relevant. Physicists’ confidence in the standard model was consolidated in 2012 when the predicted Higgs particle was eventually discovered. The extensive experimental research before and after the discovery also accumulated even more faith in the standard model. Despite the great success, there are still some concerns about the CP violation, the value of muon $g-2$ and the Higgs-mass problem. These concerns of course do not tarnish the glory of the standard model at all. But the desire for removing every single piece of doubt encourages physicists not to cease in searching for new physics.
@@ II.1 The Lagrangian and the symmetry transformations
and where now $\bar{\Psi}=\Psi^{\dagger}\Gamma^{0}$, ∂̸$\not{\partial}=\Gamma^{\mu}\partial_{\mu}∂$ with
$\begin{split}\Gamma^{0}=&\sigma^{1}\otimes I_{2}\otimes I_{2}\\ \Gamma^{k}=&i\sigma^{2}\otimes I_{2}\otimes\sigma^{k}\end{split}$ (4)
Further $\Gamma^{0}\Gamma^{0}=I_{8}$ and $\Gamma^{0}\Gamma^{k}=-\sigma^{3}\otimes I_{2}\otimes\sigma^{k}.$ Let’s assume that this Lagrangian admits the symmetry transformations of $U(4)$ among the four massless Weyl fermions:
@@ IV.1.1 The gauge configuration for deriving the Standard-Model Higgs field
Under the Lorentz transformation $\Lambda_{\frac{1}{2}m}=\exp(\theta\sigma^{3}\otimes I_{2}\otimes\sigma^{**})\cdot\exp(i\alpha I_{2}\otimes I_{2}\otimes\sigma^{1})$ with $\sigma^{**}=\cos 2\alpha\sigma^{2}-\sin 2\alpha\sigma^{3}$ $A_{mix,\mu}$ are transformed into
$\begin{split}A_{mix,0}\Rightarrow&\cosh 2\theta\phi(x)\sigma^{1}\otimes I_{2}\otimes I_{2}\\ &+i\sinh 2\theta\phi(x)\sigma^{2}\otimes I_{2}\otimes\sigma^{**},\\ A_{mix,1}\Rightarrow&-i\phi(x)\sigma^{2}\otimes I_{2}\otimes\sigma^{1},\\ A_{mix,2}\Rightarrow&i\cosh 2\theta\phi(x)\sigma^{2}\otimes\sigma^{3}\otimes\sigma^{**}\\ &-\sinh 2\theta\phi(x)\sigma^{1}\otimes\sigma^{3}\otimes I_{2},\\ A_{mix,3}\Rightarrow&-i\cos 2\alpha\phi(x)\sigma^{2}\otimes\sigma^{3}\otimes\sigma^{3}\\ &-i\sin 2\alpha\phi(x)\sigma^{2}\otimes\sigma^{3}\otimes\sigma^{2}.\end{split}$ (40)
@@
with $m_{a}+m_{b}=4\sinh^{2}\theta\cos^{2}\alpha$ and $m_{a}-m_{b}=4\sinh^{2}\theta\sin^{2}\alpha$ the mass ratios of fermions in each generation. Since $(\theta,\alpha)$ are hidden in the Gamma matrices $\Gamma_{m}^{\mu},$ they should viewed as a pair of quantum numbers. Let us calculate the Lorentz transformation explicitely
$\begin{split}\Lambda_{\frac{1}{2}m}=&\exp(\theta\sigma^{3}\otimes I_{2}\otimes\sigma^{*})\cdot\exp(i\alpha I_{2}\otimes I_{2}\otimes\sigma^{1})\\ =&\cosh\theta\cos\alpha-\sinh\theta\sigma^{3}\otimes I_{2}\otimes\sigma^{*}\\ &+i\cosh\theta\sin\alpha I_{2}\otimes I_{2}\otimes\sigma^{1}\\ \Rightarrow&\exp(\beta\sigma)\end{split}$ (43)
where $\sigma^{*}=\cos\alpha\sigma^{2}-\sin\alpha\sigma^{3},$ and the last line is added with the intention to find the generator $\sigma$ of the Lorentz transformation. But this leads to the discussion over three cases defined by the sign of the squared norm of $\sigma:$ $-\sinh^{2}\theta+\cosh^{2}\theta\sin^{2}\alpha.$ The special case where the norm is zero puts constraint on $\theta$ and $\alpha.$ Unfortunately this does not fit well with the experimental data. On the other hand, it is curious to notice that the constraint $\cosh\theta\cos\alpha=\sinh\theta$ seems to be compatible with the known information about the masses of the leptons. With this assumption and the input of the masses of electron, muon, and tauon Wiki , the masses of the neutrinos are estimated to be $\hat{m}_{\nu_{e}}=0.41ev,$ $\hat{m}_{\nu_{\mu}}=0.017Mev$ and $\hat{m}_{\nu_{\tau}}=4.92Mev.$
@@ IV.1.3 The Lagrangian of the electroweak theory—derived from a pure gauge theory
$\begin{split}L_{t}=&\sum_{m}\Psi_{m}^{\dagger}h\Gamma^{\mu}(i\partial_{\mu}+gA_{L\mu}+g^{\prime}B_{\mu})\Psi_{m}\\ &+\sum_{m}\Psi_{m}^{\dagger}g\Phi_{m}\Psi_{m}\\ &+\frac{1}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi-\frac{g^{2}}{3}[(\Phi^{\dagger}\Phi)^{2}-2\tilde{v}^{2}\Phi^{\dagger}\Phi]\\ &-\frac{1}{4}\mathbf{tr}(F_{L\mu\nu}F_{L}^{\mu\nu}+f_{Y\mu\nu}f_{Y}^{\mu\nu})\\ &+g^{2}v^{\mu}(A_{R\mu}^{3}-A_{L\mu}^{3})\Phi^{\dagger}\Phi\end{split}$ (52)
@@
with $\bar{\Phi}=i\sigma^{2}\Phi^{*}.$ The details of the calculation are presented in the appendix. An immediate observation is that there is an extra-term $g^{2}v^{\mu}(A^{3}_{L\mu}-A^{3}_{R\mu})\Phi^{\dagger}\Phi=gv^{\mu}(gA^{3}_{L\mu}-g^{\prime}p_{\mu})\Phi^{\dagger}\Phi=g\sqrt{g^{2}+g^{{}^{\prime}2}}v^{\mu}Z_{\mu}\Phi^{\dagger}\Phi$ with $v_{\mu}v^{\mu}=4/3*(246GeV)^{2}$ where $Z_{\mu}$ is the gauge field for the well-known $Z$ boson. We have not performed analysis on the quantum theory. Hopefully there are some processes that can help determine the values of $v_{\mu}.$
@@ IV.2.1 Deriving the abelian Higgs model from a gauge theory
Now just as in the non-abelian case when an internal Lorentz transformation is performed with $\Lambda_{\frac{1}{2}m}=\exp(\theta\sigma^{3}\otimes\sigma^{**})\exp(i\alpha I_{2}\otimes\sigma^{1})$ with $\sigma^{**}=\cos 2\alpha\sigma^{2}-\sin 2\alpha\sigma^{3}$, these chirality mixing gauge components are transformed into:
$\begin{split}A_{mix,0}&\Rightarrow\cosh 2\theta\phi(x)\sigma^{1}\otimes I_{2}+i\sinh 2\theta\phi(x)\sigma^{2}\otimes\sigma^{**}\\ A_{mix,1}&\Rightarrow i\phi(x)\sigma^{2}\otimes\sigma^{1}\\ A_{mix,2}&\Rightarrow i\cosh 2\theta\phi(x)\sigma^{2}\otimes\sigma^{**}-\sinh 2\theta\phi(x)\sigma^{1}\otimes I_{2}\\ A_{mix,3}&\Rightarrow i\cos 2\alpha\phi(x)\sigma^{2}\otimes\sigma^{3}+i\sin 2\alpha\sigma^{2}\otimes\sigma^{2}.\end{split}$ (60)
@@
which results in a mass ratio parameter $m=4\cosh^{2}\theta\cos^{2}\alpha.$ Alternative sign choices in $A_{mix,\mu}$ could give $m=4\sinh^{2}\theta\cos^{2}\alpha$ similar to that in the non-abelian case.
@@
$\begin{split}-\frac{1}{8}\mathbf{tr}ZF_{\mu\nu}ZF^{\mu\nu}+h.c.=&-\frac{1}{4}f_{L\mu\nu}f_{L}^{\mu\nu}-\frac{1}{4}f_{R\mu\nu}f_{R}^{\mu\nu}\\ &+\frac{1}{2}(D_{\mu}\varphi)^{\dagger}D^{\mu}\varphi\\ &-\frac{e^{2}}{3}(\varphi^{\dagger}\varphi)^{2}+\frac{e^{2}}{2}v_{\mu}v^{\mu}\varphi^{\dagger}\varphi\\ &+e^{2}v^{\mu}(b_{\mu}-a_{\mu})\varphi^{\dagger}\varphi\end{split}$ (62)
@@ V Problems and further research directions
At last, the ”over-size” of the symmetry groups in both cases raises serious concerns, as in reality only a very small fraction of the interactions described by the two full symmetry groups are observed. If some of symmetries were not to be gauged, what should be the guiding principle to discern which to be gauged and which not? One direction to look at is the effect of the transformations on the spinor metric $h.$ Under some symmetry transformations the spinor metric does not stay invariant. If the invariance of $h$ is set as a criterion for whether or not to gauge the symmetry, unfortunately transformations generated by $\sigma^{L}\otimes\sigma^{k}\otimes I_{2}$ would be excluded from the gauged symmetries. This cannot be the case. In fact, $\sigma^{L}\otimes\sigma^{k}\otimes I_{2}$ commutes with the products $h\Gamma^{\mu}$ which guarantees the products as whole entities invariant. But the products $h\Gamma^{\mu}$ would be altered by a Lorentz transformation. These arguments suggest that a simple criterion related to $h$ cannot be working. To find the true criterion, a clear understanding of the nature of the spinor metric is probably the key. Another direction is to assume that all the symmetries can be gauged, but excited with various energy-costs. After all, all the symmetries can be equally interpreted as coordinate-transformations in the internal vector space (spinors jointly formed by spin, handedness and isospin) with a norm defined. A transformation also induces a new norm. Thus making the coordinate transformation local and introduce the corresponding connection field component is a very natural thing. With the background $C_{3R}$ being postulated, excitations along some directions are going to be more energy costing. Further, it may be the case that the gauge field component associated to a non-compact transformation is more energy costing than that of a compact one. These two factors would substantially limit the number of observable interactions at the usual energy level.
@@ VI Appendices
leading to the expression of the $\Gamma^{\mu}s$ given in Eq.(4). The Pauli matrices are:
@@ VI.2 Appendix B— Calculation of the Higgs sector of the Lagrangian
$\begin{split}-\frac{1}{8}&\mathbf{tr}ZF_{\mu\nu}ZF^{\mu\nu}+h.c.=-\frac{1}{4}\{\mathbf{tr}F_{L\mu\nu}F_{L}^{\mu\nu}+\mathbf{tr}f_{R\mu\nu}f_{R}^{\mu\nu}\\ &+\mathbf{tr}f_{N\mu\nu}f_{N}^{\mu\nu}-\mathbf{tr}(D_{\mu}\eta_{\nu}-D_{\nu}\eta_{\mu})(D^{\mu}\eta^{\nu}-D^{\nu}\eta^{\mu})\\ &-g^{2}\mathbf{tr}[\eta_{\mu},\eta_{\nu}][\eta^{\mu},\eta^{\nu}]\\ &+g^{2}\mathbf{tr}([C_{3R\mu},\eta_{\nu}]-[C_{3R\nu},\eta_{\mu}])([C_{3R}^{\mu},\eta^{\nu}]-[C_{3R}^{\nu},\eta^{\mu}])\\ &-2ig\mathbf{tr}(F_{L\mu\nu}+f_{R\mu\nu}+f_{N\mu\nu})[\eta^{\mu},\eta^{\nu}]\\ &+2ig\mathbf{tr}(D_{\mu}\eta_{\nu}-D_{\nu}\eta_{\mu})([C_{3R}^{\mu},\eta^{\nu}]-[C_{3R}^{\nu},\eta^{\mu}])\}\\ &=-\frac{1}{4}\mathbf{tr}F_{L\mu\nu}F_{L}^{\mu\nu}-\frac{1}{4}\mathbf{tr}f_{R\mu\nu}f_{R}^{\mu\nu}-\frac{1}{4}\mathbf{tr}f_{N\mu\nu}f_{N}^{\mu\nu}\\ &+\frac{3}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi-3g^{2}(\Phi^{\dagger}\Phi)^{2}+\frac{3}{2}g^{2}v_{\mu}v^{\mu}\Phi^{\dagger}\Phi\\ &+3g^{2}v^{\mu}(A_{R\mu}^{3}-A_{L\mu}^{3})\Phi^{\dagger}\Phi\end{split}$ (77)
@@
Calculating the relevant terms in $-\frac{1}{4}ZF_{\mu\nu}ZF^{\mu\nu}$ gives the Higgs sector of the Lagrangian
$\begin{split}L_{\Phi}=&\frac{1}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi-\frac{g^{2}}{3}[(\Phi^{\dagger}\Phi)^{2}-2\tilde{v}^{2}\Phi^{\dagger}\Phi]\\ &+g^{2}v^{\mu}(A_{R\mu}^{3}-A_{L\mu}^{3})\Phi^{\dagger}\Phi\end{split}$ (79)
If the right-handed gauge field components $A_{R\mu}^{1,2}\sigma^{R}\sigma^{1,2}\otimes I_{2}$ are restored in the theory, the term $\frac{g^{2}}{4}\mathbf{tr}([C_{3R\mu},A_{R\nu}]-[C_{3R\nu},A_{R\mu}])([C_{3R}^{\mu},A_{R}^{\nu}]-[C_{3R}^{\nu},A_{R}^{\mu}])$ would produce two terms $-g^{2}v_{\mu}v^{\mu}A^{1}_{R\nu}A_{R}^{1\nu}+g^{2}v_{\mu}v^{\nu}A^{1}_{R\nu}A^{1\mu}_{R}$ and $-g^{2}v_{\mu}v^{\mu}A^{2}_{R\nu}A_{R}^{2\nu}+g^{2}v_{\mu}v^{\nu}A^{2}_{R\nu}A^{2\mu}_{R}$ where in the summation $\mu\neq\nu.$ When the gauge $A_{R0}^{1,2}=0$ is chosen and in the frame where $v_{k}=0$ this gives $+g^{2}v_{0}v^{0}A^{1}_{k}A^{1}_{k}$ and $+g^{2}v_{0}v^{0}A^{2}_{k}A^{2}_{k}$. Given that the vector $v_{\mu}$ must be timelike, the sign of these two terms seems to be opposite to that of a bosonic mass term in a Lagrangian. However, these terms are essentially the $E_{k}E^{k}$ part of electric-magnetic Lagrangian. Thus in the Hamiltonian, they would not change the sign and forming mass terms with correct signs. Thus, these two terms together with $\frac{1}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi$ produce six mass terms with the least mass value,
@@
While the ${W^{\pm}}$ field components would acquire a mass coefficient $\frac{1}{2}g^{2}\tilde{v}^{2}$ from $\frac{1}{2}(D_{\mu}\Phi)^{\dagger}D^{\mu}\Phi$, it can be predicted that these right-handed weak gauge bosons have masses at least $1.9$ times of the left-handed $W^{\pm}$ bosons, too heavy to be discovered at the currently running colliders. Also they are too heavy to be excited which results in the disparity between the left-handed and right-handed weak interactions.
@@
To reconstruct the abelian Higgs model, let us assume that the gauge field takes up the following configuration: $A_{\mu}=A_{L\mu}+A_{R\mu}+C_{3R\mu}+\eta_{\mu}$ with $C_{3R\mu}=v_{\mu}\sigma^{R}\otimes I_{2}$ (actually $C_{L\mu}=v_{\mu}\sigma^{L}\otimes I_{2}$ and $C_{\mu}=v_{\mu}\sigma^{3}\otimes I_{2}$ equally do the job) and
@@
$\begin{split}-\frac{1}{8}\mathbf{tr}ZF_{\mu\nu}ZF^{\mu\nu}+h.c.=&-\frac{1}{4}\mathbf{tr}f_{L\mu\nu}f_{L}^{\mu\nu}-\frac{1}{4}\mathbf{tr}f_{R\mu\nu}f_{R}^{\mu\nu}\\ &+\frac{1}{2}(D_{\mu}\varphi)^{\dagger}D^{\mu}\varphi\\ &-\frac{e^{2}}{3}(\varphi^{\dagger}\varphi)^{2}+\frac{e^{2}}{2}v_{\mu}v^{\mu}\varphi^{\dagger}\varphi\\ &+g^{2}v^{\mu}(b_{\mu}-a_{\mu})\varphi^{\dagger}\varphi\end{split}$ (82)